### v2

@@ Abstract.
[…]
@@ 2.1. Partitions
The Young diagram of a partition $\lambda$ is $Y_{\lambda}=\left\{(i,j)\in\mathbb{N}^{2}:i\leq\ell(\lambda),j\leq\lambda_{i}\right\}$ and we refer to its elements as boxes. We write $\mu\subseteq\lambda$ whenever $Y_{\mu}\subseteq Y_{\lambda}$. A skew partition $\lambda/\mu$ is a pair of partitions $\mu\subseteq\lambda$ and its Young diagram is $Y_{\lambda/\mu}=Y_{\lambda}\setminus Y_{\mu}$. We define the top of a skew partition $\lambda/\mu$, denoted as $t(\lambda/\mu)$, to be 00 if $\lambda=\mu$, and the least $i$ such that $\lambda_{i}\neq\mu_{i}$ otherwise. We similarly define the bottom $b(\lambda/\mu)$ of a skew partition by replacing the word ‘least’ with ‘greatest’.
@@
such that the skew partition $\gamma^{(i+1)}/\gamma^{(i)}$ is an $r$-border strip for all $10\leq i\leq d-$ 1 and $t(\gamma^{(1)}/\gamma^{(0)})\geq t(\gamma^{(2)}/\gamma^{(1)})\geq\dots\geq t(\gamma^{(d)}/\gamma^{(d-1)})$. If such a decomposition exists, it is unique as there is a unique choice for $\gamma^{(d-1)}$ as $\lambda/\gamma^{(d-1)}$ is an $r$-border strip with $t(\lambda/\gamma^{(d-1)})=t(\lambda/\mu)$ and an inductive argument then applies. Examples of Young diagrams of $r$-border strips and an $r$-decomposable partition are in Figure 1.
111122223333
Figure 1. Let $\gamma^{(0)}=\mu=(5,3,3,2,2,1),\gamma^{(1)}=(5,4,4,4,3,1),\gamma^{(2)}=(5,5,5,5,5,1)$ and $\gamma^{(3)}=\lambda=(8,6,6,5,5,1)$. The above dashed lines, labelled by $i=1,2,3$, pass through the Young diagrams of $5$-border strips $\gamma^{(i)}/\gamma^{(i-1)}$. The bottoms of these $5$-border strips are $5,5$ and $3$, respectively, while their tops are $2,2$ and $1$, respectively. As the tops are in non-increasing order, the skew partition $\lambda/\mu$ is $5$-decomposable.
For an $r$-border strip $\lambda/\mu$ we define its sign denoted by $\operatorname{sgn}(\lambda/\mu)$ as $(-1)^{b(\lambda/\mu)-t(\lambda/\mu)}$. For any skew partition $\lambda/\mu$ we then let $\operatorname{sgn}_{r}(\lambda/\mu)=\operatorname{sgn}(\gamma^{(1)}/\gamma^{(0)})\operatorname{sgn}(\gamma^{(2)}/\gamma^{(1)})\dots\operatorname{sgn}(\gamma^{(d)}/\gamma^{(d-1)})$ where $\gamma^{(i)}$ are as in (1) if $\lambda/\mu$ is $r$-decomposable, and $\operatorname{sgn}_{r}(\lambda/\mu)=0$ otherwise. Looking at Figure 1, the signs of the $5$-border strips there are $-1,-1$ and $1$, respectively, and hence $\operatorname{sgn}_{5}(\lambda/\mu)=1$.
@@ 2.3. Labelled abacus
Most of our terminology and notation for labelled abaci comes from [LoehrAbacus10]. A labelled abacus with $N$ beads is a sequence $w=(w_{0},w_{1},w_{2},\dots)$ indexed from 00 with precisely $N$ non-zero entries, which are $1,2,\dots,N$. For $1\leq B\leq N$, we let $w^{-1}(B)$ to be the index $i$ such that $w_{i}=B$. We write $\iota_{1}(w)>\iota_{2}(w)>\dots>\iota_{N}(w)$ for the indices $i$ such that $w_{i}$ is non-zero and define the support $\operatorname{supp}(w)$ to be $\left\{\iota_{i}(w):1\leq i\leq N\right\}$. Finally, the sign $\operatorname{sgn}(w)$ is the sign of the permutation $\sigma_{w}\in S_{N}$ given by $\sigma_{w}(B)=w_{\iota_{B}(w)}$.
@@
[…]
@@
It is easy to see how the sign changes when performing the above operations. To state the formula, we define the number of beads between positions $i_{1}<i_{2}$ as $|\operatorname{supp}(w)\cap\left\{i_{1}+1,i_{2}+2,\dots,i_{2}-1\right\}|$.
@@
In (i), $\sigma_{w^{\prime}}\sigma_{w}^{-1}$ is the transposition $(B\;C)$. In (ii), $\sigma_{w^{\prime}}\sigma_{w}^{-1}$ is a $\left(u+1\right)$-cycle. ∎
@@
111122223333444422225555111166663333$\dots$444422225555111166663333$\dots$
[…]
@@ 3. Proof of the plethystic Murnaghan–Nakayama rule
Let $w$ be the lower labelled abacus from Figure 2 and $\lambda=\operatorname{sh}(w)=(5,4,4,4,3,1)$. Since the second rightmost bead of $w$ is left-$5$-mobile, there is a corresponding $5$-border strip $\lambda/\mu$ with top $2$. Indeed, this is the $5$-border strip labelled by $1$ from Figure 1. The bijection from Lemma 3.1(iv) then pairs the labelled abaci in Figure 2.
@@
    The set $K_{N}^{r,m}(\mu,\lambda)$ is empty unless $\lambda/\mu$ is an $r$-decomposable skew partition of size $|\mu|+rm$.
@@
    If $\lambda/\mu$ is an $r$-decomposable skew partition of size $|\mu|+rm$, then the map $(w^{(0)},w^{(1)},\dots,w^{(m)})\mapsto w^{(m)}$ is a bijection from $K_{N}^{r,m}(\mu,\lambda)$ to $\operatorname{Abc}_{N}(\lambda)$.
@@
111122223333444422225555111166663333$\dots$444422225555111166663333$\dots$444422225555111166663333$\dots$444422225555111166663333$\dots$444422225555111166663333$\dots$444422225555111166663333$\dots$$(0,2,0,0,1,0)$$(0,1,0,0,1,0)$$(0,1,0,0,0,0)$$(0,0,0,0,0,0)$444422225555111166663333$\dots$444422225555111166663333$\dots$444422225555111166663333$\dots$444422225555111166663333$\dots$$(0,2,1,0,0,0)$$(0,1,1,0,0,0)$$(0,0,1,0,0,0)$$(0,0,0,0,0,0)$
Figure 3. For $N=6$, $r=5$, $m=3$ and $\mu=(5,3,3,2,2,1)$, the diagrams above the dashed line show all the values of $\alpha$ and $v$ in the process with the initial labelled abacus $w=(0,4,0,2,5,0,1,6,0,0,3,0,0,\dots)$ and the initial composition $\beta=(0,2,0,0,1,0)$. The shapes of these labelled abaci are the partitions $\gamma^{(i)}$ from Figure 1. If we use the initial composition $\beta=(0,2,1,0,0,0)$ instead, we obtain the diagrams below the dashed line. Compared to the previous diagrams, the first two $5$-moves are both with bead $2$. If we change the initial composition once more, this time to $\beta=(0,2,0,1,0,0)$, the process terminates when we reach $i=1$ as bead $4$ is not $5$-mobile.
@@
[…]
@@
The process with $(w^{\prime},\beta^{\prime})$ coincides with the process with $(w,\beta)$ where $r$-moves of bead $B$ are replaced with $r$-moves of bead $C$ as long as $\beta^{\prime}_{C}\geq(w^{\prime-1}(B)-w^{\prime-1}(C))/r$. This inequality holds true as the right-hand side is $(w^{-1}(C)-w^{-1}(B))/r$ which is at most $\beta^{\prime}_{C}$ by (6). Hence $(w^{\prime},\beta^{\prime})\in I$ and the process ends with bead $C$ $r$-colliding with bead $B$. Writing $(w^{\prime\prime},\beta^{\prime\prime})=\epsilon(w^{\prime},\beta^{\prime})$, we see that $w^{\prime\prime}=w$ and since $\operatorname{wt}_{r}(w,\beta)=\operatorname{wt}_{r}(w^{\prime},\beta^{\prime})=\operatorname{wt}_{r}(w^{\prime\prime},\beta)$, we immediately conclude that also $\beta^{\prime\prime}=\beta$.
We now move to (ii). Clearly, during the process, the weight of $(v,\alpha)$ does not change. Hence $\operatorname{wt}_{r}(w,\beta)=\operatorname{wt}_{r}(\psi(w,\beta),(0,0,\dots,0))=\operatorname{wt}(\psi(w,\beta))$, that is $\psi$ preserves weights. The rest of the claim follows from Corollary 3.4. In more detail, the statement about sizes in the ‘moreover’ part of Corollary 3.4 together with implications (ii)$\implies$(i) and (i)$\implies$(iii) shows that the map $\psi$ takes values in the desired set and is surjective, respectively. The uniqueness statement in the ‘moreover’ part shows that $\psi$ is injective and the statement about signs in the ‘moreover’ part shows that $\operatorname{sgn}(\psi(w,\beta))=\operatorname{sgn}_{r}(\lambda/\mu)\operatorname{sgn}(w,\beta)$, where $\lambda=\operatorname{sh}(\psi(w,\beta))$. ∎
@@ Acknowledgements
The author would like to thank Mark Wildon for suggesting this project and for his encouragement.
[…]

### v3

@@ Abstract.
The plethystic Murnaghan–Nakayama rule describes how to decompose the product of a Schur function and a plethysm of the form $p_{r}\circ h_{m}$ as a sum of Schur functions. We provide a short, entirely combinatorial proof of this rule using the labelled abaci introduced in [LoehrAbacus10].
@@ 2.1. Partitions
The Young diagram of a partition $\lambda$ is $Y_{\lambda}=\left\{(i,j)\in\mathbb{N}^{2}:i\leq\ell(\lambda),j\leq\lambda_{i}\right\}$ and we refer to its elements as boxes. We write $\mu\subseteq\lambda$ whenever $Y_{\mu}\subseteq Y_{\lambda}$. A skew partition $\lambda/\mu$ is a pair of partitions $\mu\subseteq\lambda$ and its Young diagram is $Y_{\lambda/\mu}=Y_{\lambda}\setminus Y_{\mu}$. We define the top of a skew partition $\lambda/\mu$, denoted as $t(\lambda/\mu)$, to be 0 if $\lambda=\mu$, and the least $i$ such that $\lambda_{i}\neq\mu_{i}$ otherwise. We similarly define the bottom $b(\lambda/\mu)$ of a skew partition by replacing the word ‘least’ with ‘greatest’.
@@
such that the skew partition $\gamma^{(i+1)}/\gamma^{(i)}$ is an $r$-border strip for all $0\leq i\leq d-1$ and $t(\gamma^{(1)}/\gamma^{(0)})\geq t(\gamma^{(2)}/\gamma^{(1)})\geq\dots\geq t(\gamma^{(d)}/\gamma^{(d-1)})$. If such a decomposition exists, it is unique as there is a unique choice for $\gamma^{(d-1)}$ as $\lambda/\gamma^{(d-1)}$ is an $r$-border strip with $t(\lambda/\gamma^{(d-1)})=t(\lambda/\mu)$ and an inductive argument then applies. Examples of Young diagrams of $r$-border strips and an $r$-decomposable skew partition are in Figure 1.
112233
Figure 1. Let $\gamma^{(0)}=\mu=(5,3,3,2,2,1),\gamma^{(1)}=(5,4,4,4,3,1),\gamma^{(2)}=(5,5,5,5,5,1)$ and $\gamma^{(3)}=\lambda=(8,6,6,5,5,1)$. The above dashed lines, labelled by $i=1,2,3$, pass through the Young diagrams of the 55-border strips $\gamma^{(i)}/\gamma^{(i-1)}$. The bottoms of these 55-border strips are $5,5$ and $3$, respectively, while their tops are $2,2$ and $1$, respectively. As the tops are in non-increasing order, the skew partition $\lambda/\mu$ is 55-decomposable.
For an $r$-border strip $\lambda/\mu$ we define its sign denoted by $\operatorname{sgn}(\lambda/\mu)$ as $(-1)^{b(\lambda/\mu)-t(\lambda/\mu)}$. For any skew partition $\lambda/\mu$ we then let $\operatorname{sgn}_{r}(\lambda/\mu)=\operatorname{sgn}(\gamma^{(1)}/\gamma^{(0)})\operatorname{sgn}(\gamma^{(2)}/\gamma^{(1)})\cdots\operatorname{sgn}(\gamma^{(d)}/\gamma^{(d-1)})$ where $\gamma^{(i)}$ are as in (1) if $\lambda/\mu$ is $r$-decomposable, and $\operatorname{sgn}_{r}(\lambda/\mu)=0$ otherwise. Looking at Figure 1, the signs of the 55-border strips there are 1,1-1,-1 and $1$, respectively, and hence $\operatorname{sgn}_{5}(\lambda/\mu)=1$.
@@ 2.3. Labelled abacus
Most of our terminology and notation for labelled abaci comes from [LoehrAbacus10]. A labelled abacus with $N$ beads is a sequence $w=(w_{0},w_{1},w_{2},\dots)$ indexed from 0 with precisely $N$ non-zero entries, which are $1,2,\dots,N$. For $1\leq B\leq N$, we let $w^{-1}(B)$ be the index $i$ such that $w_{i}=B$. We write $\iota_{1}(w)>\iota_{2}(w)>\dots>\iota_{N}(w)$ for the indices $i$ such that $w_{i}$ is non-zero and define the support $\operatorname{supp}(w)$ to be $\left\{\iota_{i}(w):1\leq i\leq N\right\}$. Finally, the sign $\operatorname{sgn}(w)$ is the sign of the permutation $\sigma_{w}\in S_{N}$ given by $\sigma_{w}(B)=w_{\iota_{B}(w)}$.
@@
If $w=(5,0,6,4,1,0,0,3,0,2,0,0,\dots)$, a labelled abacus with $6$ beads, then $\sigma_{w}=(1\;2\;3)(5\;6)$. Hence $\operatorname{sgn}(w)=-1$.
@@
It is easy to see how the sign changes when performing the above operations. To state the formula, we define the number of beads between positions $i_{1}<i_{2}$ as $|\operatorname{supp}(w)\cap\left\{i_{1}+1,i_{1}+2,\dots,i_{2}-1\right\}|$.
@@
In (i), $\sigma_{w^{\prime}}\sigma_{w}^{-1}$ is the transposition (BC)(B\;C). In (ii), $\sigma_{w^{\prime}}\sigma_{w}^{-1}$ is a $\left(u+1\right)$-cycle. ∎
@@
$\dots442255116633\dots$
Figure 2. The upper labelled abacus with $6$ beads $w$ has support $\left\{10,7,6,4,3,1\right\}$. The permutation $\sigma_{w}$ equals (1 3)(2 6 4 5)(1\;3)(2\;6\;4\;5) and thus $\operatorname{sgn}(w)=1$. We have, for instance, $w^{-1}(5)=4$ and $w^{-1}(1)=6$. The weight of $w$ is $x_{1}^{6}x_{2}^{3}x_{3}^{10}x_{4}x_{5}^{4}x_{6}^{7}$ and the shape of $w$ is $(5,3,3,2,2,1)$. Bead $4$ is not $5$-mobile, but the other beads are. By 55-moving bead $2$ we obtain the lower labelled abacus $w^{\prime}$. One computes that $\sigma_{w^{\prime}}\sigma_{w}^{-1}=(2\;5\;1\;6)$, which is in accordance with the proof of Lemma 2.3(ii).
@@ 3. Proof of the plethystic Murnaghan–Nakayama rule
Let $w$ be the lower labelled abacus from Figure 2 and $\lambda=\operatorname{sh}(w)=(5,4,4,4,3,1)$. Since the second rightmost bead of $w$ is left-55-mobile, there is a corresponding 55-border strip $\lambda/\mu$ with top 22. Indeed, this is the 55-border strip labelled by 11 from Figure 1. The bijection from Lemma 3.1(iv) then pairs the labelled abaci in Figure 2.
@@
    The set $K_{N}^{r,m}(\mu,\lambda)$ is empty unless $\lambda/\mu$ is an $r$-decomposable skew partition of size $rm$.
@@
    If $\lambda/\mu$ is an $r$-decomposable skew partition of size $rm$, then the map $(w^{(0)},w^{(1)},\dots,w^{(m)})\mapsto w^{(m)}$ is a bijection from $K_{N}^{r,m}(\mu,\lambda)$ to $\operatorname{Abc}_{N}(\lambda)$.
@@
[…]
Figure 3. For $N=6$, $r=5$, $m=3$ and $\mu=(5,3,3,2,2,1)$, the diagrams above the dashed line show all the values of $\alpha$ and $v$ in the process with the initial labelled abacus $w=(0,4,0,2,5,0,1,6,0,0,3,0,0,\dots)$ and the initial composition $\beta=(0,2,0,0,1,0)$. The shapes of these labelled abaci are the partitions $\gamma^{(i)}$ from Figure 1. If we use the initial composition $\beta=(0,2,1,0,0,0)$ instead, we obtain the diagrams below the dashed line. Compared to the previous diagrams, the first two 55-moves are both with bead $2$. If we change the initial composition once more, this time to $\beta=(0,2,0,1,0,0)$, the process terminates when we reach $i=1$ as bead $4$ is not 55-mobile.
@@
$\beta^{\prime}_{j}=\begin{cases*}\beta_{B}-\frac{w^{-1}(C)-w^{-1}(B)}{r}&if $j=B$,\\ \beta_{C}+\frac{w^{-1}(C)-w^{-1}(B)}{r}&if $j=C$,\\ \beta_{j}&otherwise.\end{cases*}$ (6)
@@
The process with $(w^{\prime},\beta^{\prime})$ coincides with the process with $(w,\beta)$ where $r$-moves of bead $B$ are replaced with $r$-moves of bead $C$ as long as $\beta^{\prime}_{C}\geq(w^{\prime-1}(B)-w^{\prime-1}(C))/r$. This inequality holds true as the right-hand side is $(w^{-1}(C)-w^{-1}(B))/r$ which is at most $\beta^{\prime}_{C}$ by (6). Hence $(w^{\prime},\beta^{\prime})\in I$ and the process ends with bead $C$ $r$-colliding with bead $B$. Writing $(w^{\prime\prime},\beta^{\prime\prime})=\epsilon(w^{\prime},\beta^{\prime})$, we see that $w^{\prime\prime}=w$ and since $\operatorname{wt}_{r}(w,\beta)=\operatorname{wt}_{r}(w^{\prime},\beta^{\prime})=\operatorname{wt}_{r}(w^{\prime\prime},\beta^{\prime\prime})$, we immediately conclude that also $\beta^{\prime\prime}=\beta$.
[…]
@@ Acknowledgements
The author would like to thank Mark Wildon for suggesting this project and for his encouragement, as well as the anonymous referees for their comments, which helped greatly improve the manuscript.